OpenStudy (emma.davenport):

can someone help me simplify this?? (a^3b^2c/a^5c^2)(a^2bc^4)

OpenStudy (anonymous):

\[\frac{a^3 b^2 c \left(a^2 b c^4\right)}{a^5 c^2}=b^3 c^3 \]

OpenStudy (emma.davenport):

\[((a^3b^2c)/a^5c^2)*(a^2b/c)\]